Solve for the unknown interest rate in each of the following: Complete the following analysis. Do not hard code values in youAll answers must be entered as formula.
Answer #1
Future Value = Present Value*((1+r)^t)
where r is the interest rate and t is the time period in years.
1) 297 = 240*((1+r)^4)
(1+r)^4 = 1.2375
1+r = (1.2375)^(1/4)
1+r = 1.054718
r = .054718
r = 5.47%.
2) 1080 = 360*((1+r)^18)
(1+r)^18 = 1080/360
(1+r)^18 = 3
1+r = (3)^(1/18)
1+r = 1.062935
r = .062935
r = 6.29%.
3) 185382 = 39000*((1+r)^19)
(1+r)^19 = 185382/39000
(1+r)^19 = 4.753385
1+r = (4.753385)^(1/19)
1+r = 1.085505
r = .085505
r = 8.55%.
4) 531618 = 38261*((1+r)^25)
(1+r)^25 = 531618/38261
(1+r)^25 = 13.89451
1+r = (13.89451)^(1/25)
1+r = 1.110999
r = .110999
r = 11.09%.